Related Rates: Problem Solving


Direction: Solve each word problem involving a related rates. Make a sketch of your work

if necessary, and then apply differentiation.


Out of frustration, Arceli throws a stone into a river. This caused ripples in the form

of concentric circles. The radius of the outer ripple increases constantly at 1.5 feet

per second. At what rate is the total area of the disturbed water changing when the

radius is 5 feet?


PLEASE ANSWER MY QUESTION QUICKLY!!

DEADLINE : 05/03/2022 11 : 00 PM



1
Expert's answer
2022-05-05T17:13:40-0400

"A=\pi r^2"

"dA/dt=\pi 2r(dr/dt)"

r=5

dr/dt=1.5

"dA/dt=\pi 5(1.5)2=47.1" sq. feet per second.